# What Scales in Cross-Entropy Scaling Law?

### v1

@@ 1 Introduction
However, the cross-entropy scaling law has recently faced growing skepticism, both in practice and in theory. Practitioners have raised concerns about whether the law can continue to provide accurate predictions on larger scales. Empirical studies show that while cross-entropy decreases with an accurate power-law trend for small models, this trend becomes noticeably slower for very large models (Kaplan et al., 2020; Chen et al., 2025b). On the theoretical side, the situation is equally unsettled. Although most existing theoretical frameworks can somehow prove the scaling law for error-based metrics, such as mean squared error (Lyu et al., 2025), they cannot directly generalize to cross-entropy loss. Thus, the theoretical foundation of this law is still unclear. These problems have fueled skepticism about whether cross-entropy truly scales, and in turn, they undermine confidence in scaling up models as a reliable path forward.
In this work, we hypothesize that it is not cross-entropy loss itself that truly scales, but rather a dominant component hidden within it. This component gives the illusion that the cross-entropy itself follows a scaling law. Identifying this scaling component would therefore be of significant value. On the practical side, it would provide a more reliable law to guide the development of large language models. On the theoretical side, it could offer a better objective for investigating the fundamental principles of artificial intelligence. Motivated by this, the research question of this paper is:
@@
Building on this decomposition, we investigate the scaling behavior of these components and find that only Error-Entropy truly scales. We refer to it as the Error-Entropy Scaling Law. Concretely, we conduct experiments on multiple real-world language datasets using thirty-two models spanning five orders of magnitude in size.  [^1] The code is available at https://anonymous.4open.science/r/RethinkCE-9357/. We then evaluate the scaling behavior of cross-entropy alongside its three decomposed components. The results are clear. (1) Error-Entropy decreases according to a power law with respect to model size, whereas the other two terms remain random for different model sizes. (2) The power-law fit of Error-Entropy is even better than that of cross-entropy. It indicates that the Error-Entropy scaling law is potentially the reason why cross-entropy approximately exhibits scaling behavior.
We believe that our decomposition of cross-entropy and the discovery of the Error-Entropy scaling law provide a better description of how language models actually scale, and open the door to broad applications. As one example, this framework helps resolve a long-standing puzzle (Kaplan et al., 2020; Chen et al., 2025b): why does the cross-entropy scaling law appear accurate for small models but break down for larger ones? The answer lies in the proportion of Error-Entropy. It dominates in small models, making cross-entropy follow a clean power law, but its declining share in larger models allows non-scaling terms to dominate, breaking the law. This specific example highlights how Error-Entropy can sharpen our understanding of model behavior. More broadly, we expect that it will have wide-ranging implications, such as guiding the design and training of large language models, probing model mechanisms, and finding fundamental theories of artificial intelligence.
@@ 2 Related Work
Cross-Entropy Loss. In the neural network domain, prior studies about cross-entropy have mainly examined its properties regarding consistency, robustness, calibration, and regularization (Guo et al., 2017; Fort & Scherlis, 2018; Wei et al., 2022b; Soudry et al., 2024; Vysogorets et al., 2024). For instance, Mao et al. (2023) conducted a systematic theoretical analysis of cross-entropy and its variants, deriving H-consistency bounds. However, these analyses do not investigate whether—and how—such intrinsic properties of cross-entropy interact with macroscopic scaling behavior. Our work complements these directions by decomposing cross-entropy into components with clearer operational meaning and studying how these components scale with model size.
@@ 3.1 Rank-based Error
Therefore, we propose a new metric, Rank-based Error (RBE), which equals the rank of the ground-truth token predicted by the language model. It serves as a measure of how far the prediction deviates from the right answer. Formally, let $i$ be the index of the input data, and $v_{i}$ be the correct token for this data. $\mathcal{V}$ is the vocabulary. We use $s_{v_{i}}$ to denote the model’s output probability score for $v_{i}$. Thus, RBE for $v_{i}$ is defined as:
@@
Based on RBE, we define two distributions. First, we treat RBE as a random variable and analyze its probability distribution over the corpus. Second, we group model predictions based on RBE values and compute the mean output scores for each group. An illustration is shown in Fig. 2. The details are elaborated as follows:
@@ 3.2 Mathematical Decomposition
This section mathematically decomposes cross-entropy. We start with its definition: cross-entropy measures the divergence between the predicted distribution and the ground-truth distribution. In language modeling, the ground-truth distribution is usually implemented as a one-hot vector that takes the value $1$ at the ground-truth token and 0 elsewhere. Let $s_{v_{i}}$ be the output score for ground-truth token $v_{i}$. Then cross-entropy loss is $\log s_{v_{i}}$. Given a corpus, cross-entropy loss is the average of all $\log s_{v_{i}}$.
@@ 4 Error-Entropy Scaling Law
4.1 Qualitative Analysis
In this section, we study how the three components of our cross-entropy decomposition—Error-Entropy (EE), Self-Alignment (SA), and Confidence (Conf)—scale with model size. [^3] Experimental details for this section are provided in App. A.2. The results are shown in Fig. 7.
@@ 4.2 Quantitative Analysis
Now we further examine the scaling properties of the decomposed components with quantitative fitting experiments.
We use two metrics, namely $R^{2}$ to examine how well the results follow a power-law rate, and $\Delta$ to examine how close the scaling behavior is compared to the scaling of cross-entropy. They are defined as follows. Let $N$ be the number of model non-embedding parameters. For each metric $M\in\{\mathrm{CE},\mathrm{EE},\mathrm{SA},\mathrm{Conf}\}$, we perform a log-log linear regression across models and report the goodness of fit $R^{2}$:
@@ 5 Application of Error-Entropy Scaling Law
Our decomposed result and the discovered Error-Entropy scaling law better characterize the model scaling behavior. We believe they will have wide applications for large language models. In this section, we provide an example in using them to understand the puzzle of slowing cross-entropy scaling law.
The puzzle is that cross-entropy scales precisely at a power-law rate for small models, and yet slows down for very large models. Concretely, for small models, Kaplan et al. (2020) described the relationship between cross-entropy $\mathcal{L}_{\mathrm{CE}}$ and model scale $M$ as $\mathcal{L}_{\mathrm{CE}}\propto M^{-\alpha}$, which is a pure power-law form. Yet, when model size grows, Chen et al. (2025a) found cross-entropy loss starts to slow down. To account for this problem, the expression is rewritten as $\mathcal{L}_{\mathrm{CE}}\propto M^{-\alpha}+bias$. Nowadays, researchers even start to find out cross-entropy slows down even more (Qin et al., 2025; Lourie et al., 2025). This phenomenon has been widely observed but poorly understood.
@@
This is just one application example. Since our findings provide a fundamental understanding of model behavior, we believe they have wide-ranging applications. For example, our findings also suggest new directions for training objectives, such as retaining only the Error-Entropy term as training loss. We leave these explorations to future work.
@@ 6 Conclusion
Appendix A More Details of Experimental Setups
@@ A.1 Experimental Setup for Training Dynamics
Unless otherwise noted, we evaluate on the first 1 000 instances from each dataset (after shuffling with a fixed seed of 42). We further verified robustness by repeating selected experiments with substantially larger sample sizes (e.g., 10k instances), and observed identical trends, indicating that our conclusions are insensitive to the particular sample count or ordering.
@@ A.2 Experimental Setup for Scaling
Data. We separately sample 1 000 English instances from Wikipedia (20231101 snapshot) (Guo et al., 2020), C4 (Dodge et al., 2021), and the GitHub portion of The Pile (Gao et al., 2020). All instances are drawn with a fixed random seed 42 from the training split and reused across models to ensure comparability.

### v2

@@ 1 Introduction
However, the cross-entropy scaling law has recently faced growing skepticism, both in practice and in theory. Practitioners have raised concerns about whether the law can continue to provide accurate predictions on larger scales. Empirical studies show that while cross-entropy decreases with an accurate power-law trend for small models, this trend becomes noticeably slower for very large models (Kaplan et al., 2020; Chen et al., 2025). On the theoretical side, the situation is equally unsettled. Although most existing theoretical frameworks can somehow prove the scaling law for error-based metrics, such as mean squared error (Lyu et al., 2025), they cannot directly generalize to cross-entropy loss. Thus, the theoretical foundation of this law is still unclear. These problems have fueled skepticism about whether cross-entropy truly scales, and in turn, they undermine confidence in scaling up models as a reliable path forward.
In this work, we hypothesize that it is not cross-entropy loss itself that truly scales, but rather a dominant component hidden within it. This component gives the illusion that the cross-entropy itself follows a scaling law. Identifying this scaling component would therefore be of significant value. On the practical side, it would provide a more reliable law to guide the development of large language models. On the theoretical side, it could offer a better objective for investigating the principles of artificial intelligence. Motivated by this, the research question of this paper is:
@@
Building on this decomposition, we investigate the scaling behavior of these components and find that only Error-Entropy truly scales. We refer to it as the Error-Entropy Scaling Law. Concretely, we conduct experiments on multiple real-world language datasets using 32 models spanning five orders of magnitude in size.  [^1] The code is available at https://github.com/yanjx2021/RethinkCE. We then evaluate the scaling behavior of cross-entropy alongside its three components. The results are clear. (1) Error-Entropy decreases according to a power law with model size, whereas the other two terms remain random for different model sizes. (2) The power-law fit of Error-Entropy is even better than that of cross-entropy. It indicates that the Error-Entropy scaling law is potentially the reason why cross-entropy approximately exhibits scaling behavior.
We believe that our decomposition of cross-entropy and the discovery of the Error-Entropy scaling law provide a better description of how language models actually scale, and open the door to broad applications. As one example, this framework helps resolve a long-standing puzzle (Kaplan et al., 2020; Chen et al., 2025): why does the cross-entropy scaling law appear accurate for small models but break down for larger ones? The answer lies in the proportion of Error-Entropy. It dominates in small models, making cross-entropy follow a clean power law, but its declining share in larger models allows non-scaling terms to dominate, breaking the law. This specific example highlights how Error-Entropy can sharpen our understanding of model behavior. More broadly, we expect that it will have wide-ranging implications, such as guiding the design and training of large language models, probing model mechanisms, and finding fundamental theories of artificial intelligence.
@@ 2 Related Work
Cross-Entropy Loss. In the neural network domain, prior studies about cross-entropy have mainly examined its properties regarding consistency, robustness, calibration, and regularization (Guo et al., 2017; Fort and Scherlis, 2018; Wei et al., 2022b; Soudry et al., 2024; Vysogorets et al., 2024). For instance, Mao et al. (2023) conducted a systematic theoretical analysis of cross-entropy and its variants, deriving H-consistency bounds. However, these analyses do not investigate whether—and how—such intrinsic properties of cross-entropy interact with macroscopic scaling behavior. Our work complements these directions by decomposing cross-entropy into components with clearer operational meaning and studying how these components scale with model size.
@@ 3.1 Rank-based Error
Therefore, we propose a new metric, Rank-based Error (RBE), which equals the rank of the ground-truth token predicted by the language model. It serves as a measure of how far the prediction deviates from the right answer. Formally, let $i$ be the index of the input data, and $v_{i}$ be the correct token for this data. $\mathcal{V}$ is the vocabulary. We use $s_{v_{i}}$ to denote the model’s output probability score for $v_{i}$.
Thus, RBE for $v_{i}$ is defined as:
@@
Based on RBE, we define two distributions. First, we treat RBE as a random variable and analyze its probability distribution over the corpus. Second, we group model predictions by RBE and compute the mean score for each group. An illustration is shown in Fig. 2. The details are as follows:
@@ 3.2 Mathematical Decomposition
This section mathematically decomposes cross-entropy. We start with its definition: cross-entropy measures the divergence between the predicted distribution and the ground-truth distribution. In language modeling, the ground-truth distribution is usually implemented as a one-hot vector that takes the value $1$ at the ground-truth token and 0 elsewhere. Let $s_{v_{i}}$ denote the score of ground-truth token $v_{i}$. Then cross-entropy loss is $\log s_{v_{i}}$. Given a corpus, it is the average of all $\log s_{v_{i}}$.
@@ 4 Error-Entropy Scaling Law
4.1 Qualitative Analysis
In this section, we study how the three decomposed components—Error-Entropy (EE), Self-Alignment (SA), and Confidence (Conf), scale with model size. [^3] Experimental details for this section are provided in App. A.2. The results are shown in Fig. 7.
@@ 4.2 Quantitative Analysis
We further examine the scaling properties of the components with quantitative fitting experiments.
We use two metrics, namely $R^{2}$ to examine how well the results follow a power-law rate, and $\Delta$ to examine how close the scaling behavior is to that of cross-entropy. They are defined as follows. Let $N$ be the number of model non-embedding parameters. For each metric $M\in\{\mathrm{CE},\mathrm{EE},\mathrm{SA},\mathrm{Conf}\}$, we perform a log-log linear regression across models and report the goodness of fit $R^{2}$:
@@ 5 Application of Error-Entropy Scaling Law
Our decomposed result and the discovered Error-Entropy scaling law better characterize the model scaling behavior. We believe they will have wide applications for LLMs. In this section, we provide an example in using them to understand the puzzle of slowing cross-entropy scaling law.
The puzzle is that cross-entropy scales precisely at a power-law rate for small models, and yet slows down for very large models. Concretely, for small models, Kaplan et al. (2020) described the relationship between cross-entropy $\mathcal{L}_{\mathrm{CE}}$ and model scale $M$ as $\mathcal{L}_{\mathrm{CE}}\propto M^{-\alpha}$, which is a pure power-law form. Yet, when model size grows, OpenAI et al. (2024) found cross-entropy loss starts to slow down. To account for this problem, the expression is rewritten as $\mathcal{L}_{\mathrm{CE}}\propto M^{-\alpha}+bias$. Nowadays, researchers even start to find out cross-entropy slows down even more (Qin et al., 2025; Lourie et al., 2025). This phenomenon has been widely observed but poorly understood.
@@
6 Disscusion and Future Work
In this section, we further discuss the contributions and theoretical implications of our work. We analyze the deeper principles behind Error-Entropy in Sec. 6.1. We also outline possible training objectives and future directions motivated by this perspective in Sec. 6.2.
6.1 Deeper principles behind Error-Entropy
Although our definition of Error-Entropy (EE) arises from cross-entropy decomposition, we believe it reflects deeper underlying principles. In information-theoretic learning (ITL) and signal processing field, error entropy has been studied extensively (Erdogmus and Principe, 2002; Liu et al., 2006; Li et al., 2022). Our work establishes a link between these fields and large language models, offering a new perspective on model behavior and training dynamics.
While our formulation of Error-Entropy differs from the classic ITL version, the two are closely connected. ITL studies focus on nonparametric regression, where both input and output are real-valued. The error is defined as $e=y-\hat{y_{pred}}$, which makes optimizing the error distribution a natural objective. A common ITL-style error entropy is: $H(e)=-\int p(e)\log p(e)\,de.$ In contrast, language modeling is a classification task where each token in the vocabulary is assigned a probability, and the goal is to raise the rank of the correct token. We define a rank-based error that restores an error distribution in this setting, and use it to decompose cross-entropy into error-related and probability-related components. Our error-entropy shares a similar form with the ITL definition, which links the probability-based formulation used in LLM training to the error-distribution perspective from ITL.
This connection paves the way for a new line of research. From an optimization perspective, techniques developed in ITL kernel-based methods for minimizing error entropy (Chen et al., 2013; 2019) may inspire new training objectives for language models. These approaches often use stochastic gradient methods in kernel spaces, and have been shown to handle heavy-tailed and non-Gaussian noise effectively. From a theoretical perspective, existing insights from ITL can be revisited in the LLMs setting to explain the observed scaling behavior of EE. In particular, the robustness of error-entropy under noisy or mismatched supervision, one of its original motivations in ITL (Tang and Li, 2011), may also help explain why EE scales more consistently than CE in large-scale models.
6.2 Training objectives inspired by Error-Entropy
In this subsection, we show that it is feasible to design practical training objectives inspired by Error-Entropy. EE itself is non-differentiable with respect to the logits, so we instead consider surrogate training objectives. Specifically, we define a loss that adds a penalty term on Confidence:
$\mathcal{L}_{\lambda}=\text{CE}+\lambda\cdot\text{CONF},\quad 0<\lambda<1.$ (10)
We penalize the Confidence term because it is not directly related to ranking ability. As discussed in Section 3.3, increasing Confidence alone does not bring significant improvement in performance. Moreover, our empirical results (Fig. 8) shows that its share in cross-entropy grows with model size, which suggests that current training over-emphasizes this component. By reducing its weight in the loss, we aim to shift optimization toward EE, which is the component with more desired scaling.
Mathematically, the compensated loss is differentiable with respect to the logits, so it can be optimized with standard gradient-based training. In our notation, cross-entropy can be written as $\text{CE}=-\frac{1}{N}\sum_{i}\log s_{i}$, where $s_{i}$ is the score of the correct token for example $i$. The Confidence term takes the form $\text{CONF}=\log\sum_{e}Q_{e}$, where $Q_{e}$ is the mean score over examples with error level $e$. Differentiating $\mathcal{L}_{\lambda}$ with respect to $s_{i}$ yields (See App. D for the full derivation.)
$\frac{\partial\mathcal{L}_{\lambda}}{\partial s_{i}}=-\frac{1}{Ns_{i}}\Big(1-\lambda\frac{q_{e}}{p_{e}}\Big),$ (11)
where $p_{e}$ is the error distribution and $q_{e}$ is the score distribution. With original cross-entropy loss, once a token attains the correct rank, further increasing its probability only raises $q_{e}$ without improving $p_{e}$. The factor $\bigl(1-\lambda\frac{q_{e}}{p_{e}}\bigr)$ encourages the model to keep $q_{e}$ closer to $p_{e}$, so that optimization focuses more on improving the error distribution captured by error-entropy.
Beyond such differentiable surrogates, error-entropy can also be used directly as a non-differentiable objective in reinforcement-learning style fine-tuning, where it serves as a reward signal that depends only on ranks. This avoids differentiating error-entropy directly and may provide a complementary route to align training with error-entropy.
@@ 7 Conclusion
Appendix A More Details of Experimental Setup
@@ A.1 Experimental Setup for Training Dynamics
Unless otherwise noted, we evaluate on the first 1 000 instances from each dataset (after shuffling with a fixed seed of 42). Note that these instances are full sequences containing many tokens; the total number of evaluation tokens is summarized in Table 3. We further verified robustness by repeating selected experiments with substantially larger sample sizes (e.g., 10k instances), and observed identical trends, indicating that our conclusions are insensitive to the particular sample count or ordering.
@@
Table 3: Total number of evaluation tokens used for the main experiments.
Wikipedia C4 Pile-GitHub
# Tokens 704,185 417,140 1,091,481
@@ A.2 Experimental Setup for Scaling
Datasets.
Sampling protocol.
Unless otherwise noted, we evaluate on the first 1 000 instances from each dataset (after shuffling with a fixed seed of 42). Note that these instances are full sequences containing many tokens; the total number of evaluation tokens is summarized in Table 3. We further verified robustness by repeating selected experiments with substantially larger sample sizes (e.g., 10k instances), and observed identical trends, indicating that our conclusions are insensitive to the particular sample count or ordering.
@@ A.2.2 Models for Quantitative Scaling Analysis
Appendix B Extensive Training Dynamics Experiments
In Figure 9, we illustrate how cross-entropy and its decomposed components evolve throughout training. The plots cover three datasets and track training progress on Pythia-12B, one of the largest open-source models with available checkpoints. This confirms our main results by showing that all three components are indeed optimized during training.
Figure 9: Evolution of cross-entropy and its components during training across three datasets with Pythia-12B.
Appendix C Extensive Qualitative Analysis
In Figure 10, we report extended qualitative results for cross-entropy and its decomposed components across model scales. This figure uses the same set of 32 models as in our main quantitative analysis and further includes Meta-Llama-3.1-405B, one of the largest open-source language models currently available.
The overall trends are fully consistent with our main results. Both cross-entropy and Error-Entropy exhibit clear scaling behavior, with Error-Entropy lying systematically closer to the fitted power-law line. In contrast, the Self-Alignment and Confidence terms do not show a stable decreasing pattern with model size and display much larger variance. Importantly, adding the 405B-parameter model does not change these qualitative conclusions, reinforcing our claim that Error-Entropy is the robust component follows a scaling law in the open-model regime we can study.
Appendix D Derivation of the compensated loss gradient
We collect here the notation and give a detailed derivation of the gradient for the compensated loss
$\mathcal{L}_{\lambda}=\mathrm{CE}+\lambda\cdot\mathrm{CONF},\quad 0<\lambda<1.$ (12)
Notation.
We consider a corpus with $N$ evaluation tokens indexed by $i=1,\dots,N$. For each token $i$: $s_{i}$ denotes the softmax score of the ground-truth token, $e_{i}$ denotes its rank-based error level.
For each error level $e$, we group tokens via $I_{e}=\{i:e_{i}=e\},n_{e}=|I_{e}|$, and define the error distribution $p_{e}=n_{e}/N$ and the geometric mean score $Q_{e}=\bigl(\prod_{i\in I_{e}}s_{i}\bigr)^{1/n_{e}}.$ We then aggregate over error levels by $C=\sum_{e}Q_{e}$ and $q_{e}=Q_{e}/C,$
In this notation, the corpus-level cross-entropy and Confidence terms are
$\mathrm{CE}$ $=-\frac{1}{N}\sum_{i=1}^{N}\log s_{i},$ (13)
$\mathrm{CONF}$ $=\log C=\log\sum_{e}Q_{e}.$ (14)
Derivation.
We treat the scores $\{s_{i}\}$ as independent variables.
Cross-entropy term. The derivative of cross-entropy with respect to $s_{i}$ is
$\frac{\partial\mathrm{CE}}{\partial s_{i}}=-\frac{1}{N}\cdot\frac{1}{s_{i}}=-\frac{1}{Ns_{i}}.$ (15)
Confidence term. For the geometric mean $Q_{e}$ we have
$\log Q_{e}=\frac{1}{n_{e}}\sum_{j\in I_{e}}\log s_{j}.$ (16)
Thus, for a token $i$ with error level $e=e_{i}$,
$\frac{\partial\log Q_{e}}{\partial s_{i}}=\frac{1}{n_{e}}\cdot\frac{1}{s_{i}},\qquad\Rightarrow\qquad\frac{\partial Q_{e}}{\partial s_{i}}=Q_{e}\cdot\frac{1}{n_{e}s_{i}}.$ (17)
Since $C=\sum_{e}Q_{e}$, only the term at $e=e_{i}$ depends on $s_{i}$, so
$\frac{\partial C}{\partial s_{i}}=\frac{\partial Q_{e_{i}}}{\partial s_{i}}=\frac{Q_{e_{i}}}{n_{e_{i}}s_{i}}.$ (18)
Using $\mathrm{CONF}=\log C$, we obtain
$\frac{\partial\mathrm{CONF}}{\partial s_{i}}=\frac{1}{C}\cdot\frac{\partial C}{\partial s_{i}}=\frac{Q_{e_{i}}}{C\,n_{e_{i}}s_{i}}=\frac{q_{e_{i}}}{n_{e_{i}}s_{i}},$ (19)
where we used $q_{e}=Q_{e}/C$. Recalling that $p_{e}=n_{e}/N$ so that $n_{e}=Np_{e}$, we can rewrite this as
$\frac{\partial\mathrm{CONF}}{\partial s_{i}}=\frac{q_{e_{i}}}{Np_{e_{i}}s_{i}}=\frac{1}{Ns_{i}}\cdot\frac{q_{e_{i}}}{p_{e_{i}}}.$ (20)
Compensated loss. Combining equation 15 and equation 20, the derivative of $\mathcal{L}_{\lambda}=\mathrm{CE}+\lambda\cdot\mathrm{CONF}$ with respect to $s_{i}$ is
$\frac{\partial\mathcal{L}_{\lambda}}{\partial s_{i}}$ $=\frac{\partial\mathrm{CE}}{\partial s_{i}}+\lambda\frac{\partial\mathrm{CONF}}{\partial s_{i}}$ (21)
$=-\frac{1}{Ns_{i}}+\lambda\cdot\frac{1}{Ns_{i}}\cdot\frac{q_{e_{i}}}{p_{e_{i}}}$ (22)
$=-\frac{1}{Ns_{i}}\Bigl(1-\lambda\frac{q_{e_{i}}}{p_{e_{i}}}\Bigr).$ (23)
Finally, since each score $s_{i}$ is a smooth function of the logits via the softmax transformation, the compensated loss $\mathcal{L}_{\lambda}$ is differentiable with respect to the logits, and can be optimized with standard gradient-based training.
Figure 10: Extended illustration of cross-entropy and its decomposed components over model size.
Appendix E Statement on the Use of Large Language Models
We used large language models (LLMs) solely for polishing the writing of this manuscript, including grammar correction, style harmonization, and minor sentence-level rewrites. All technical content—definitions, theorems, methods, experiments, and conclusions—was authored and verified by the authors. Any LLM-suggested phrasing that was adopted was manually reviewed and edited. No unpublished data, proprietary code, or sensitive information was provided to any LLM. This usage does not affect the reproducibility of our methods or results. This disclosure follows the ICLR 2026 policy on LLM usage.